OpenStudy (renren):

Can I get some help with this problem. Thanks The mass of planet mercury is 3.3*10^23 kilograms. The mass of planet Jupiter is 5.8*10^3 times the mass of planet Mercury. Planet Jupiter is 3.2*10^2 times the mass of planet Earth. Part A Using scientific notation, what is the mass of planet Jupiter? Answer in Kilogram Part B Using scientific notation, what is the mass of planet Earth? Answer in Kilogram

OpenStudy (cutebunny12):

its B

OpenStudy (anonymous):

for jupiter, it is (3.3x5.8)x10^23x10^2 (When multiplying exponents w/ same base, add the powers. ex: 10^2 x 10^4 = 10^6) to find the mass of earth, divide the mass of jupiter by 3.2x10^2.

OpenStudy (cutebunny12):

ok

OpenStudy (renren):

Can U give me the answer for A and B. Thanks

OpenStudy (renren):

For A I get 1.914.10^26

OpenStudy (anonymous):

correct. 1.914x10^26 is correct for A

OpenStudy (renren):

Part B I get 3.2*10^2=320 thats the answer I gets

OpenStudy (anonymous):

B is 1.914x10^26 (jupiters mass) / 3.2x10^2 1.914/3.2 = 0.0598125 --> 5.98125x10^24

OpenStudy (renren):

Thank you again Robjoe

OpenStudy (anonymous):

@renren no prob!

OpenStudy (renren):

Ur welcome